# ESOD: Efficient Small Object Detection on High-Resolution Images

### v1

@@ Abstract
Enlarging input images is a straightforward and effective approach to promote small object detection. However, simple image enlargement is significantly expensive on both computations and GPU memory. In fact, small objects are usually sparsely distributed and locally clustered. Therefore, massive feature extraction computations are wasted on the non-target background area of images. Recent works have tried to pick out target-containing regions using an extra network and perform conventional object detection, but the newly introduced computation limits their final performance. In this paper, we propose to reuse the detector’s backbone to conduct feature-level object-seeking and patch-slicing, which can avoid redundant feature extraction and reduce the computation cost. Incorporating with a sparse detection head, we are able to detect small objects on high-resolution inputs (e.g., 1080P or larger) for superior performance. The resulting Efficient Small Object Detection (ESOD) approach is a generic framework, which can be applied to both CNN- and ViT-based detectors to save the computation and GPU memory costs. Extensive experiments demonstrate the efficacy and efficiency of our method. In particular, our method consistently surpasses the SOTA detectors by a large margin (e.g., 8% gains on AP) on the representative VisDrone, UAVDT, and TinyPerson datasets. Code will be made public soon.
@@ I Introduction
With recent advances in convolutional neural networks (CNNs) [1, 2, 3] and Vision Transformers (ViTs) [4, 5, 6], general object detection has achieved promising performance on public benchmarks including MS COCO [7, 8] and Pascal VOC [9, 10]. It has become the foundation for widespread applications, such as autonomous driving [11, 12] and security monitoring [13, 14]. However, detecting small objects (e.g., less than $32\times 32$ pixels [7] remains a challenging task [15, 16, 17], which hinds the visual analysis on specialized platforms like unmanned aerial vehicles (UAVs) [18, 19] and panoramic cameras [13, 20].
@@
    We statistically state the fact of sparsely clustered small objects in practice, and conduct feature-level object-seeking with adaptive patch-slicing to avoid redundant feature extraction and save the computation cost. Moreover, sparse detection head is adopted to reuse the estimated objectness mask for further computation-saving.
@@
    We propose a generic framework ESOD, which can adapt to both CNN and ViT architectures, so as to save the computation and GPU memory when detecting high-resolution images.
@@ II-A Small Object Detection
Compared with the above works, input image enlarging [27, 43] is more effective but lacks efficiency. Our work focuses on exploring a way to save computation and speed up detection when enlarging images.

### v2

@@ Abstract
Enlarging input images is a straightforward and effective approach to promote small object detection. However, simple image enlargement is significantly expensive on both computations and GPU memory. In fact, small objects are usually sparsely distributed and locally clustered. Therefore, massive feature extraction computations are wasted on the non-target background area of images. Recent works have tried to pick out target-containing regions using an extra network and perform conventional object detection, but the newly introduced computation limits their final performance. In this paper, we propose to reuse the detector’s backbone to conduct feature-level object-seeking and patch-slicing, which can avoid redundant feature extraction and reduce the computation cost. Incorporating with a sparse detection head, we are able to detect small objects on high-resolution inputs (e.g., 1080P or larger) for superior performance. The resulting Efficient Small Object Detection (ESOD) approach is a generic framework, which can be applied to both CNN- and ViT-based detectors to save the computation and GPU memory costs. Extensive experiments demonstrate the efficacy and efficiency of our method. In particular, our method consistently surpasses the SOTA detectors by a large margin (e.g., 8% gains on AP) on the representative VisDrone, UAVDT, and TinyPerson datasets. Code is available at https://github.com/alibaba/esod.
@@ I Introduction
With recent advances in convolutional neural networks (CNNs) [1, 2, 3] and Vision Transformers (ViTs) [4, 5, 6], general object detection has promising achieves on public benchmarks including MS COCO [7, 8] and Pascal VOC [9, 10]. It has become the foundation for widespread applications, such as autonomous driving [11, 12] and security monitoring [13, 14]. However, detecting small objects (e.g., less than $32\times 32$ pixels [7]) remains a challenge [15, 16, 17], which hinders visual analysis on specialized platforms like unmanned aerial vehicles (UAVs) [18, 19] and panoramic cameras [13, 20].
@@
    We statistically state the fact of sparsely clustered small objects in practice, and conduct feature-level object-seeking with adaptive patch-slicing to avoid redundant feature extraction and save the computation cost. A sparse detection head is adopted to reuse the estimated objectness mask for further computation-saving.
@@
    We propose a generic framework ESOD that adapts to both CNN and ViT architectures, saving detection computation and GPU memory on high-resolution images.
@@ II-A Small Object Detection
In contrast, input image enlarging [27, 42] is more effective but lacks efficiency, and our work focuses on saving computation and speeding up detection when enlarging images.
@@ V Conclusion
Acknowledgment
This work was supported in part by the Fundamental Research Funds for the Central Universities, in part by Alibaba Cloud through the Research Intern Program, and in part by Zhejiang Provincial Natural Science Foundation of China under Grant No. LDT23F01013F01. The work of Kai Liu was completed when he was with Alibaba Cloud.